Bob is riding his bike (combined mass of 90 kg) over a hill, which creates a radius of 30m. He is traveling at a speed of 4 m/s. What is the force he feels from his bike as he is going over the crest of the hill?


1
Expert's answer
2021-05-30T13:20:56-0400

"F=F_1+F_2=mg+\frac{mv^2}{r}=m(g+\frac{v^2}{r}),"

"F=90\cdot (9.8+\frac{4^2}{30})=930~\text N."